I'm confused by this problem: "Find the x-intercepts y = x^2 + 3x". Can anyone show me how to solve and find the x-intercepts on this type of equation? All of the examples in my text involve equations with constant terms (something like y = x^2 + 3x + 6). This one has me a bit confused. Thanks.
Answer by ankor@dixie-net.com(22740)

I'm confused by this problem: "Find the x-intercepts y = x^2 + 3x". Can anyone show me how to solve and find the x-intercepts on this type of equation? All of the examples in my text involve equations with constant terms (something like y = x^2 + 3x + 6). This one has me a bit confused. Thanks.
:
It's even easier
:
y = 0
x^2 + 3x = 0
:
Factor out x:
x(x + 3) = 0
therefore:
x = 0
and
x = -3
:
The x intercepts are 0 and -3